OpenStudy (anonymous):

Trying to figure out the problem with my daughter probability. a bag contains 3 red marbles and 7 blue marbles. a marble is drawn and replaced. use an equation to show the probability of drawing 3 red marbles

OpenStudy (anonymous):

You want the probability of drawing 3 times, replacing them each time, and getting a red on each draw? 3/10 * 3/10 * 3/10 = 27/1000 or 0.027

OpenStudy (doc.brown):

I won't take that bet.

OpenStudy (anonymous):

Yup that is the correct response

OpenStudy (doc.brown):

In case you were wondering, if you kept the red ones out: 3/10 * 2/9 * 1/8 = 0.0004

OpenStudy (anonymous):

Love probability questions. Got anymore for us?

OpenStudy (anonymous):

Basically the probability of picking a red marble is \[\frac{\text{red marbles}}{\text{ total amount of marbles }}=\frac{3}{10}\] Now since we keep replacing the marble once picked back into the bag the probability of picking a red marble is always 3/10 for each drawing Hence we get (3/10)*(3/10)*(3/10)

OpenStudy (anonymous):

@jnyberg-homework Do you follow?

OpenStudy (anonymous):

Sorry it wasn't working on my phone she has to have an equation...the title of the lesson heading its calculating frequency distribution in saxon math

OpenStudy (anonymous):

well the equation is (3/10)*(3/10)*(3/10)=27/1000

OpenStudy (anonymous):

I think variables have to be involved

OpenStudy (anonymous):

I don't think you are giving us the complete question. To write an equation, something must be unknown or vary.

OpenStudy (anonymous):

Yaa im pretty stuck myself. I dont see anything varying

OpenStudy (anonymous):

Is the number of draws to get 3 reds unknown?

OpenStudy (anonymous):

it is 3 red marbles in a row...notice that was missing

OpenStudy (anonymous):

Hmmm any examples in the book to demonstrate what you are referring to

OpenStudy (anonymous):

I wish there were this particular book is not the best. Up until this book they were great...

OpenStudy (anonymous):

I think they mean that they dont want you to just give an answer they want you to show how you reached your answer. What equation did you use and so the equation you used was (3/10)*(3/10)*(3/10)

OpenStudy (anonymous):

glad it wasn't we both went huh

OpenStudy (anonymous):

@BangkokGarrett This seems like a very basic probability class in HS. I think its pretty straight forward and should NOT include variables

OpenStudy (anonymous):

8th grade algebra 1

OpenStudy (anonymous):

She had probability early int he year & this is experimental probability

OpenStudy (anonymous):

dunno to me this seems extremely basic and would not include any variables Showing an equation does not mean it must include variables. It means show the set up of your equation and how you arrived at the solution.

OpenStudy (anonymous):

I think I will have her put that & we will see tomorrow what they were after... I do have one other question she is asking for help on, do I start a new question?

OpenStudy (anonymous):

You can ask it here and if I can be of assistance I will help :)

OpenStudy (anonymous):

a square has area 9m^4-54m^3+81m^2 square inches what is its side length?

OpenStudy (anonymous):

GCF= Greatest Common Factor

OpenStudy (anonymous):

Fine First factor out the GCF from all 3 terms...which is 9m^2 9m^4-54m^3+81m^2 9m^2 * (m^2 - 6m + 9) 9m^2 * (m - 3)^2 (3m)^2 * (m - 3)^2 That's = [3m(m - 3)]^2 So each side of the square is 3m(m - 3) or 3m^2 - 9m

OpenStudy (anonymous):

@jnyberg-homework Do you follow the solution?

OpenStudy (anonymous):

Sorry we had a quick issue, looking at it now

OpenStudy (anonymous):

Which of the answers do you think would be better?

OpenStudy (anonymous):

What do you mean? He only provided one answer

OpenStudy (anonymous):

the square version or the 3m(m-3) is what my daughter meant